A HeNe laser has an HR mirror with a reflectivity of 99.9%. The transmission of the output coupler is 1.0%, and the combined scattering and absorption loss at the mirror surface is 0.04%. The saturated amplifier gain is 1.0073. Find the roundtrip cavity loss.
Answer #1

The reflectivity of the HR mirror is given to be 99.9 % = 0.999

The output coupler transmission is 1%

The scattering and absorption loss at this mirror is 0.04%

So, total loss = 1.04%

So, the reflectivity of mirror = 100-1.04 = 98.96% = 0.9896

The amplifier gain = 1.0073

So,

The round trip cavity loss is given by

L = R1*R2*G (assuming that there is no transmission loss in the medium)

Using the values,

L = 0.999*0.9896*1.0073 = 0.9958

The value is less than 1. This means that the cavity loses energy in one round trip (not good for laser)